# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

### v1

@@ 2.4. Weak div stability
We now prove the following important commutative property.
@@
Notice that $\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$. So
$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}I_{k}^{\rm tn}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
@@
For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$. Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (8) implies the desired result (9). ∎
If the velocity space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$. Since DoF (6b) to define $I_{k}^{\rm tn}$ is for moments of degree k1k-1, the commutative property (8) no longer holds. Consequently, the inf-sup condition may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
@@
Since $\operatorname{div}_{w}\Sigma_{k}^{\rm tn}=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, the operator $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ is injective, and thus $\|\operatorname{dev}{\rm grad\,}_{w}(\cdot)\|$ defines a norm on $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
The inf-sup condition (9) can then be written equivalently as
@@
Consider the mixed finite element method for the vector Laplacian
$\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{ with }\;\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d}).$
Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ such that
(11) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
The formulation (11) is well posed by the inf-sup conditions (9) or (10). However, the errors $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled. As a result, the dominant contribution is the $H^{1}$-type error of $\boldsymbol{u}$ of order $h^{k}$, rather than the $L^{2}$ error of $\boldsymbol{\sigma}$ of order $h^{k+1}$. This suboptimality comes from the loss of exact divergence-free conformity:
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\ \ \forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\ \ \nRightarrow\ \ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},E\boldsymbol{v})_{0,h}=0\ \ \forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
@@ 2.5. Commuting property for the weak dev grad operator
As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]|_{F}\in\mathbb{P}_{k}(F)$, we have
$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F},$
$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F}$
@@
The weak operator $\operatorname{div}_{w}$ can be regarded as a discretization of $\operatorname{div}$ in the distributional sense. We can preserve the exact divergence condition in the following sense.
@@
then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$.
@@ 2.6. Restriction to divergence-free subspaces
The conditions (8) and (14) are incompatible. Weak div stability holds only on the smaller space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, while the exact divergence-freeness requires the larger space $\mathbb{V}_{k,k}^{\operatorname{div}}$. We resolve this conflict by considering the divergence-free subspaces.
We have the following commuting property (cf. [5]):
(15) $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=Q_{\ell}\big(\operatorname{div}\boldsymbol{v}\big),\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}),$
which implies $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=0$ for all $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$. The additional interior moments in ${\rm RT}_{k}$ compared with ${\rm BDM}_{k}$ serve only to enrich the range of the divergence operator. Hence, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ and $k\geq 1$,
$I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
@@
Let $k\geq 0$. The mixed finite element method (17) is well posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (16). Then
@@
The well-posedness follows from the inf-sup condition (9) and the fact
Denote by $\boldsymbol{\sigma}_{I}=I_{k}^{\rm tn}\boldsymbol{\sigma}$ and $\boldsymbol{u}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. By (8), it holds that $(\operatorname{div}_{w}\boldsymbol{\sigma}_{I},\boldsymbol{v}_{h})_{0,h}=(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v}_{h})=-(\boldsymbol{f},\boldsymbol{v}_{h})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$, so we conclude
@@
We verify the following orthogonality by using (14)
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}$
$=(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=-(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{I})_{0,h}=0.$
@@
The velocity space $\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ and $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, and the optimal order of convergence for $H^{1}$ type error is $h^{k}$. So (20) is a superconvergent error estimate for $\boldsymbol{u}$.
3. Mixed Finite Element Methods for Stokes Equation
In this section, we introduce the pressure as a Lagrange multiplier to enforce the divergence-free condition, and another multiplier to relax the tangential-normal continuity of the stress space. The resulting mixed finite element methods for the Stokes equation are pointwise divergence-free and superconvergent.

### v2

@@ 2.4. Weak div stability
We now prove the following important commutative property which can be simply written as
$Q_{k,k-1}^{\operatorname{div}}\operatorname{div}=\operatorname{div}_{w}I_{k}^{\rm tn},$
where $Q_{k,k-1}^{\operatorname{div}}$ is the $L^{2}$-projection to the space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
@@
Notice that $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\!]|_{F}\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$ for $F\in\mathcal{F}_{h}$. So
$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(I_{k}^{\rm tn}\boldsymbol{\tau})\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}_{h}]\!])_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}_{h}]\!])_{F}$
@@
For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, the surjectivity of the divergence operator mapping from $H^{1}(\Omega;\mathbb{T})$ onto $L^{2}(\Omega;\mathbb{R}^{d})$ guarantees that there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$ and $\|\boldsymbol{\tau}\|_{1}\lesssim\|\boldsymbol{v}_{h}\|$ (cf. [6, (33)]). Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (12) implies $\operatorname{div}_{w}\boldsymbol{\tau}_{h}=\boldsymbol{v}_{h}$, which, combined with the boundedness of the interpolant $I_{k}^{\rm tn}$, yields the desired result (13). ∎
Remark 2.6.
If the range space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$ not $\Sigma_{k-1}^{-1}(\mathbb{T})$. Since DoF (10b) to define $I_{k}^{\rm tn}$ is for moments of degree k1k-1, the commutative property (12) no longer holds. Consequently, the inf-sup condition (13) may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
Remark 2.7.
On the other hand, to accommodate the slip boundary condition, we can reduce the stress space to impose the boundary condition $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}=\boldsymbol{0}$ on $\partial\Omega$. Let
$\mathring{H}^{1}(\Omega;\mathbb{T}):=\{\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T}):\Pi_{F}\boldsymbol{\tau}\boldsymbol{n}=\boldsymbol{0}\text{ on }\partial\Omega\}.$
Following the proof of (13), and using the surjectivity $\operatorname{div}\mathring{H}^{1}(\Omega;\mathbb{T})=L^{2}(\Omega;\mathbb{R}^{d})$ together with the commuting property (12), one can show that
$\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\mathring{\Sigma}_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha$
for some constant $\alpha>0$ independent of $h$, where
$\mathring{\Sigma}_{k}^{\rm tn}:=\{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}:\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}=\boldsymbol{0}\text{ on }\partial\Omega\}.$
@@
Since $\operatorname{div}_{w}\Sigma_{k}^{\rm tn}=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, the operator $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ is injective, and thus $\|\operatorname{dev}{\rm grad\,}_{w}(\cdot)\|$ defines a norm on $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$. Consequently, the inf-sup condition (13) can be written equivalently as
@@
With these preparations, we consider the mixed finite element method for the vector Laplacian: Find $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ such that $\Delta\boldsymbol{u}=\boldsymbol{f}$. The discrete problem seeks $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ satisfying:
(15) $\begin{cases}(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=0&\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},\\ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=-(\boldsymbol{f},\boldsymbol{v}_{h})&\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.\end{cases}$
The mixed method (15) is well-posed due to the inf-sup conditions (13) and (14). However, a standard error analysis reveals that $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled. Subtracting the continuous system from (15) yields:
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\quad\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
$(\operatorname{div}_{w}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}),\boldsymbol{v}_{h})_{0,h}$ $=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
The term $(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=(\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h},\boldsymbol{\tau}_{h})$ is bounded by the $H^{1}$-type error $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|$ rather than by $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ alone. This suboptimality arises because being weakly divergence-free with respect to $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ does not imply strong divergence-free properties:
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\quad\nRightarrow\quad\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=0\quad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
However, we will demonstrate in Lemma 2.9 that this property is preserved if the velocity test space in (15) is enriched from $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. In this enriched setting, the test space is sufficiently large to ensure that for $\boldsymbol{\tau}_{h}=I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}$, $\operatorname{div}_{w}\boldsymbol{\tau}_{h}=0$ holds. Consequently, the coupling term $(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=0$ vanishes, yielding the optimal stress estimate. By decoupling the stress error from the $O(h^{k})$ velocity approximation error, we avoid the lower-order contamination and achieve the full $O(h^{k+1})$ convergence rate.
@@ 2.5. Commuting property for the weak dev grad operator
As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!]|_{F}\in\mathbb{P}_{k}(F)$, we have
$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\![\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!])_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\![\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!])_{F}$
@@
The weak operator $\operatorname{div}_{w}$ can be regarded as a discretization of $\operatorname{div}$ in the distributional sense. We can preserve the strong divergence-free property in the following sense.
@@
then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$ in distribution sense.
@@ 2.6. Restriction to divergence-free subspaces
The commuting conditions (12) and (18) involve different velocity spaces, as shown in diagram (19):
(19) $H^{1}(\Omega;\mathbb{T})$$L^{2}(\Omega;\mathbb{R}^{d})$$\Sigma_{k}^{\rm tn}$$\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$$\scriptstyle{\operatorname{div}}$$\scriptstyle{I_{k}^{\rm tn}}$$\scriptstyle{Q_{k,k-1}^{\operatorname{div}}}$$\scriptstyle{\operatorname{div}_{w}}$$H^{1}(\Omega;\mathbb{R}^{d})$$L^{2}(\Omega;\mathbb{T})$$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$$\Sigma_{k}^{\rm tn}$$\scriptstyle{\operatorname{dev}{\rm grad\,}}$$\scriptstyle{I_{k,k}^{\operatorname{div}}}$$\scriptstyle{Q_{k}^{\rm tn}}$$\scriptstyle{\operatorname{dev}{\rm grad\,}_{w}}$
Weak divergence stability holds on the smaller space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, whereas strong divergence-freeness requires the larger space $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. We resolve this structural mismatch by restricting our analysis to the divergence-free subspace.
Recall the standard property for $H(\operatorname{div})$ interpolants [5]:
(20) $\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})=Q_{\ell}(\operatorname{div}\boldsymbol{v}),\quad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
Equation (20) implies that $I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\in\ker(\operatorname{div})$ whenever $\boldsymbol{v}\in\ker(\operatorname{div})$. Notably, for $k\geq 1$, the additional interior moments in $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$ compared to $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ only enrich the range of the divergence operator and do not alter the divergence-free part of the space. Consequently, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$, we have:
(21) $I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
This identity bridges the two diagrams, allowing us to leverage the stability of $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ alongside the consistency of the weak deviatoric gradient on $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
@@
By the inf-sup condition (13) and the commuting property (18), we obtain the following two properties.
Lemma 2.10.
There exists a constant $\alpha>0$, independent of $h$, such that
(24) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha.$
Proof.
It follows from the inf-sup condition (13) and the fact
Lemma 2.11.
We have the strong divergence-free property: if $(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}=0$ for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$, then in the distributional sense:
$\langle\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle=0\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
Proof.
By the definition of distributional $\operatorname{div}$, the commuting property (18), and (21)
$\langle\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle=-(\boldsymbol{\tau}_{h},\operatorname{dev}{\rm grad\,}\boldsymbol{v})=-(\boldsymbol{\tau}_{h},\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{v})=-(\operatorname{div}_{w}\boldsymbol{\tau}_{h},I_{k,k}^{\operatorname{div}}\boldsymbol{v})=0.$
@@
Let $k\geq 0$. The mixed finite element method (23) is well-posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (22). Then
@@
The well-posedness is from (24). Denote by $\boldsymbol{\sigma}_{I}=I_{k}^{\rm tn}\boldsymbol{\sigma}$ and $\boldsymbol{u}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. By (12), it holds that $(\operatorname{div}_{w}\boldsymbol{\sigma}_{I},\boldsymbol{v}_{h})_{0,h}=(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v}_{h})=-(\boldsymbol{f},\boldsymbol{v}_{h})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$, so we conclude
@@
So by Lemma 2.11, $(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=0$. We obtain the following orthogonality
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}=0.$
@@
Since $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, the standard approximation theory for the velocity space $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ yields an $H^{1}$-type error of order $O(h^{k})$. Consequently, the estimate (27), which establishes a higher convergence rate $O(h^{k+1})$, constitutes a velocity superconvergence result.
3. Mixed Finite Element Methods for the Stokes Equation
In this section, we develop a mixed finite element discretization for the Stokes equation that ensures the pointwise divergence-free velocity. We introduce Lagrange multiplier spaces using piecewise polynomials for pressure and facial multipliers to relax the tangential-normal continuity of the pseudo-stress. After proving stability via a discrete inf-sup condition, we formulate the mixed scheme and show its equivalence to (23). The resulting error analysis yields optimal-order, pressure-robust bounds. Finally, we establish $L^{2}$-velocity superconvergence of order $O(h^{k+2})$ and present a postprocessing strategy that achieves $H^{1}$-superconvergence.